# sangkuriang: A pseudo-spectral Python library for Korteweg-de Vries soliton simulation

### v2

@@ 2.3 Numerical Experiments
The sangkuriang solver implementation is validated and demonstrated through four progressively complex test cases that probe different aspects of KdV soliton dynamics. These scenarios are designed to assess: (i) baseline propagation characteristics of isolated solitons, (ii) interaction dynamics between equal-amplitude waves, (iii) overtaking collision phenomena arising from amplitude-dependent velocities, and (iv) complex multi-body interactions in three-soliton systems. While the sangkuriang framework is general and may be applied to arbitrary initial conditions satisfying the periodicity constraint, these canonical test cases provide standardized benchmarks that have been extensively studied in the literature [2, 6, 3] and serve to verify the implementation’s ability to capture the essential physics of nonlinear dispersive wave propagation.
The first test case examines the propagation of a single soliton in isolation, serving as a baseline reference for subsequent multi-soliton scenarios. The initial condition is specified using the exact single-soliton solution profile (33) in the form
@@
The second test case investigates the interaction between two solitons of equal amplitude, which may exhibit phase shift phenomena without energy exchange. The initial condition comprises a linear superposition of two sech2 profiles:
@@
The third test case examines the overtaking collision between solitons of unequal amplitudes, demonstrating the remarkable elastic scattering property that defines soliton behavior. The initial condition takes the form of equation (52) with asymmetric parameters: $A_{1}=6.0$ m, $w_{1}=1.5$ m, $x_{1}=-18.0$ m for the leading (taller) soliton, and $A_{2}=2.0$ m, $w_{2}=2.5$ m, $x_{2}=-5.0$ m for the trailing (shorter) soliton. According to the velocity relation (34), the taller soliton propagates at $v_{1}=0.4$ m/s while the shorter moves at $v_{2}\approx 0.133$ m/s, ensuring that the faster wave overtakes the slower one during the simulation. The domain spans $x\in[-40,40]$ m with $N=512$ points, and the final time is extended to $T_{\text{final}}=70.0$ s to observe the complete collision sequence and subsequent separation. This scenario probes the nonlinear interaction dynamics most severely, as the collision involves significant amplitude modulation and temporary waveform distortion [2]. The remarkable theoretical prediction—that both solitons emerge from the collision with their original amplitudes and velocities intact, differing only by phase shifts—provides a stringent test of the numerical method’s ability to capture the integrable structure of the KdV equation [5]. Conservation errors may increase during the collision phase due to enhanced nonlinearity, but should remain below $10^{-4}$ for the employed tolerances.
The fourth test case extends the complexity to a three-soliton system, creating a choreography of multiple interactions that tests the sangkuriang solver’s robustness under more challenging conditions. The initial condition generalizes equation (52) to three components:
@@
with amplitudes $A_{1}=7.0$ m, $A_{2}=4.0$ m, $A_{3}=2.5$ m, widths $w_{1}=1.2$ m, $w_{2}=1.8$ m, $w_{3}=2.2$ m, and initial positions $x_{1}=-25.0$ m, $x_{2}=-10.0$ m, $x_{3}=5.0$ m, respectively. The corresponding velocities are $v_{1}\approx 0.467$ m/s, $v_{2}\approx 0.267$ m/s, and $v_{3}\approx 0.167$ m/s, ordered such that each faster soliton will sequentially overtake the slower ones ahead. To accommodate the increased spatial extent of the three-wave system and provide adequate resolution for the narrowest soliton, the domain is enlarged to $x\in[-50,50]$ m with enhanced resolution $N=1024$ ($\Delta x\approx 0.098$ m). The simulation extends to $T_{\text{final}}=80.0$ s with tightened tolerances $\text{rtol}=10^{-11}$ and $\text{atol}=10^{-13}$ to maintain accuracy through multiple collision events. This test case may exhibit complex interaction patterns including transient amplitude amplifications during collisions, the generation of small-amplitude dispersive radiation, and intricate phase shift accumulations [6, 3]. The successful navigation of these multiple interactions while preserving conservation laws and soliton identities demonstrates the sangkuriang solver’s capability for research-grade simulations of multi-soliton dynamics.
@@
The physical parameters $\mu=0.1$ m3/s and $\varepsilon=0.2$ m-1 are held constant across all test cases to facilitate direct comparison of dynamical behaviors. These values are representative of shallow water wave systems and ensure that the balance between nonlinearity and dispersion falls within the regime where soliton solutions are well-defined [26]. The specific choices of soliton amplitudes and positions in each test case are informed by established benchmarks in the soliton literature [2, 20] and are designed to produce qualitatively distinct interaction scenarios within computationally tractable simulation times. While these four cases provide comprehensive validation of the sangkuriang solver’s core functionality, the implementation supports arbitrary initial conditions through the SechProfile, TanhProfile, GaussianProfile, and MultiSoliton classes, enabling users to investigate alternative scenarios such as soliton generation from non-sech2 initial data, wave packet decomposition, or soliton interactions with continuous wave backgrounds.
@@
The output from each sangkuriang simulation includes the complete spatiotemporal field $u(x,t)$, conservation law time series $M(t)$, $P(t)$, $E(t)$ as defined in equations (45)–(47), relative conservation errors (48)–(50), and diagnostic information regarding computational cost (number of function evaluations, wall clock time). Data are stored in netCDF4 format [23] following CF-1.8 metadata conventions, facilitating interoperability with standard scientific data analysis tools. Animated visualizations are generated showing the evolving wave profile in three-dimensional space-time representations, with conservation law diagnostics overlaid to provide immediate visual feedback on solution quality. These outputs enable both qualitative assessment of soliton dynamics and quantitative verification of the numerical implementation’s fidelity to the underlying physics encoded in the KdV equation (29).
@@ 2.4 Data Analysis
The diagnostic analysis of simulation outputs is performed independently of the sangkuriang solver library through a dedicated suite of post-processing routines implemented in Python. These routines employ NumPy [11] for array operations and FFT, SciPy [12] for signal processing and distance computations, and Matplotlib [35] for visualization. Simulation data archived in netCDF4 format [23] following CF-1.8 conventions are subjected to four complementary analyses: verification of conservation law preservation, soliton trajectory tracking with velocity validation, spectral information-theoretic characterization, and phase space recurrence quantification. Each analysis extracts quantitative metrics that collectively assess both numerical accuracy and physical fidelity of the computed solutions.
@@
provides a time-averaged measure of conservation fidelity that is less sensitive to isolated outliers. To distinguish systematic drift from bounded oscillations, the linear drift rate $\gamma_{Q}$ is extracted via ordinary least-squares regression. Defining the design matrix $\mathbf{A}\in\mathbb{R}^{N_{t}\times 2}$ with elements $A_{n1}=t_{n}$ and $A_{n2}=1$, and the observation vector $\mathbf{b}=[\delta_{Q}(t_{0}),\delta_{Q}(t_{1}),\ldots,\delta_{Q}(t_{N_{t}-1})]^{T}$, the drift rate is obtained as the first component of the least-squares solution
@@
Soliton trajectory extraction proceeds through systematic identification and tracking of local amplitude maxima in the spatiotemporal field $u(x_{j},t_{n})$. At each output time $t_{n}$, the discrete spatial profile $\{u_{j}^{(n)}\}_{j=0}^{N-1}\equiv\{u(x_{j},t_{n})\}_{j=0}^{N-1}$ is analyzed using the scipy.signal.find_peaks routine [12], which identifies indices $j^{*}$ satisfying $u_{j^{*}}^{(n)}>u_{j^{*}-1}^{(n)}$, $u_{j^{*}}^{(n)}>u_{j^{*}+1}^{(n)}$, subject to additional constraints. Two filtering parameters govern peak selection: a height threshold $u_{\text{th}}$ requiring $u_{j^{*}}^{(n)}>u_{\text{th}}$, and a minimum separation distance $d_{\text{min}}$ (in grid points) enforcing $|j_{1}^{*}-j_{2}^{*}|>d_{\text{min}}$ between distinct peaks. For the test cases examined, $u_{\text{th}}=0.5$ m for multi-soliton configurations and $u_{\text{th}}=1.0$ m for single-soliton runs, with $d_{\text{min}}=10$ grid points throughout. Each detected peak yields a triplet $(t_{n},x_{j^{*}},u_{j^{*}}^{(n)})$ recording the time, spatial position, and amplitude.
For configurations containing $K$ solitons, individual trajectories must be disambiguated from the aggregate peak data. This is accomplished through amplitude-based sorting at each time instant: letting $\{(x_{k}^{(n)},A_{k}^{(n)})\}_{k=1}^{K_{n}}$ denote the $K_{n}$ peaks detected at time $t_{n}$ with positions $x_{k}^{(n)}$ and amplitudes $A_{k}^{(n)}$, these are reordered such that $A_{1}^{(n)}\geq A_{2}^{(n)}\geq\cdots\geq A_{K_{n}}^{(n)}$. The $k$-th soliton track is then defined as the sequence $\{(t_{n},x_{k}^{(n)})\}$ across all times where at least $k$ peaks were detected. This ranking scheme exploits the amplitude-velocity ordering inherent to KdV solitons: since taller solitons propagate faster according to equation (34), amplitude rank provides a consistent identifier even through collision events where spatial ordering temporarily inverts.
@@
where terms with $p_{m}^{(n)}<10^{-15}$ are excluded from the summation to avoid numerical singularities. The normalization by $\ln N$ ensures $S_{k}^{(n)}\in[0,1]$, with $S_{k}^{(n)}=0$ corresponding to a monochromatic signal (all energy concentrated at a single wavenumber) and $S_{k}^{(n)}=1$ to uniform spectral distribution (white noise). Soliton profiles, characterized by smooth $\text{sech}^{2}$ envelopes with exponentially decaying Fourier spectra, exhibit intermediate entropy values that remain approximately constant during propagation.
@@
measures the squared Euclidean distance between the observed distribution $\{p_{m}^{(n)}\}$ and the uniform distribution $\{1/N\}$. This construction ensures $C_{\text{LMC}}^{(n)}=0$ for both perfectly ordered states ($H^{(n)}=0$, all energy at one mode) and maximally disordered states ($D^{(n)}=0$, uniform distribution), attaining nonzero values only for intermediate configurations exhibiting nontrivial structure [38]. The LMC complexity thus distinguishes structured nonequilibrium states from both trivial order and featureless randomness.
@@
with one-sided differences at domain boundaries. Higher Fisher information indicates sharper spatial gradients and hence more localized wave structures; for soliton solutions, $F^{(n)}$ scales inversely with the characteristic width parameter $w$ appearing in equation (33). During soliton collisions, transient increases in $F^{(n)}$ reflect the temporary steepening of gradients as wave profiles overlap.
@@
measures the fraction of recurrence points forming diagonal line structures of length at least $\ell_{\min}$ (typically $\ell_{\min}=2$). Here $P(\ell)$ denotes the number of diagonal lines of exactly length $\ell$ in the recurrence matrix, where a diagonal line at offset $k>0$ consists of consecutive entries $R_{i,i+k}=R_{i+1,i+1+k}=\cdots=1$. Diagonal structures arise when segments of the trajectory evolve in parallel, a hallmark of deterministic dynamics on regular (non-chaotic) attractors [43]. For integrable systems such as the KdV equation, DET $\to 1$ reflects confinement to invariant tori, whereas chaotic systems exhibit DET $\ll 1$ due to sensitive dependence on initial conditions disrupting parallel trajectory segments.
@@
3 Results and Discussion
The sangkuriang solver was validated through four test cases of increasing complexity, executed on a laptop-class workstation (Lenovo ThinkPad P52s with Intel Core i7-8550U processor at 4.0 GHz, 8 logical cores) running Fedora Linux 39 (kernel 6.11.9-100.fc39.x86_64). This hardware configuration represents a modest computational platform, permitting assessment of the solver’s practical utility for research applications without access to high-performance computing resources.
Table 1 summarizes the computational cost for each test case. The single-soliton simulation (Case 1) required 208,202 adaptive time steps to integrate over the 50 s interval, completing in approximately 152 s of wall-clock time. The adaptive time stepper selected finer temporal resolution for this case due to the smaller spatial domain ($L=60$ m) and correspondingly higher wavenumber content relative to Cases 2 and 3. The two equal-amplitude soliton simulation (Case 2) completed in 75.7 s using 98,612 steps, while the overtaking collision (Case 3) required 115,658 steps and 87.6 s. The three-soliton configuration (Case 4), employing doubled spatial resolution ($N=1024$) and tightened tolerances ($\text{rtol}=10^{-11}$, $\text{atol}=10^{-13}$), demanded 551,858 steps and 534 s of computation time. These timings suggest that the Numba-accelerated implementation achieves reasonable throughput for exploratory research, though the scaling with grid resolution and tolerance warrants consideration for production simulations.

### v3

@@ 2.3 Numerical Experiments
The sangkuriang solver implementation is validated and demonstrated through four progressively complex test cases that probe different aspects of KdV soliton dynamics relevant to idealized ocean wave propagation. These scenarios are designed to assess: (i) baseline propagation characteristics of isolated solitary waves, as observed when a single internal soliton traverses a continental shelf [46]; (ii) interaction dynamics between equal-amplitude waves, analogous to tidally generated soliton trains with similar amplitudes [49]; (iii) overtaking collision phenomena arising from amplitude-dependent velocities, a common occurrence when faster, larger-amplitude internal waves catch up with smaller predecessors in a wave packet [47, 51]; and (iv) complex multi-body interactions in three-soliton systems, representing the rich dynamics of rank-ordered internal soliton trains frequently observed in satellite imagery of marginal seas [48]. While the sangkuriang framework is general and may be applied to arbitrary initial conditions satisfying the periodicity constraint, these canonical test cases provide standardized benchmarks that have been extensively studied in the literature [3, 6, 1] and serve to verify the implementation’s ability to capture the essential physics of nonlinear dispersive wave propagation in shallow water environments.
The first test case examines the propagation of a single soliton in isolation, serving as a baseline reference for subsequent multi-soliton scenarios. This configuration represents the simplest idealization of a solitary internal wave propagating across a flat-bottomed shelf region. The initial condition is specified using the exact single-soliton solution profile (33) in the form
@@
The second test case investigates the interaction between two solitons of equal amplitude, which may exhibit phase shift phenomena without energy exchange. This scenario is motivated by observations of tidally generated internal wave packets in which successive solitons emerge with comparable amplitudes from a generation site such as a sill or shelf break [49, 51]. The initial condition comprises a linear superposition of two sech2 profiles:
@@
The third test case examines the overtaking collision between solitons of unequal amplitudes, demonstrating the remarkable elastic scattering property that defines soliton behavior. In the ocean, such overtaking events occur routinely when a train of rank-ordered internal solitary waves propagates across a shelf, with the leading (largest) soliton moving fastest according to the amplitude–velocity relation (34), and occasionally overtaking smaller waves that were generated earlier or at a different source [47, 48]. The initial condition takes the form of equation (52) with asymmetric parameters: $A_{1}=6.0$ m, $w_{1}=1.5$ m, $x_{1}=-18.0$ m for the leading (taller) soliton, and $A_{2}=2.0$ m, $w_{2}=2.5$ m, $x_{2}=-5.0$ m for the trailing (shorter) soliton. According to the velocity relation (34), the taller soliton propagates at $v_{1}=0.4$ m/s while the shorter moves at $v_{2}\approx 0.133$ m/s, ensuring that the faster wave overtakes the slower one during the simulation. The domain spans $x\in[-40,40]$ m with $N=512$ points, and the final time is extended to $T_{\text{final}}=70.0$ s to observe the complete collision sequence and subsequent separation. This scenario probes the nonlinear interaction dynamics most severely, as the collision involves significant amplitude modulation and temporary waveform distortion [3]. The remarkable theoretical prediction—that both solitons emerge from the collision with their original amplitudes and velocities intact, differing only by phase shifts—provides a stringent test of the numerical method’s ability to capture the integrable structure of the KdV equation [5]. Conservation errors may increase during the collision phase due to enhanced nonlinearity, but should remain below $10^{-4}$ for the employed tolerances.
The fourth test case extends the complexity to a three-soliton system, creating a choreography of multiple interactions that tests the sangkuriang solver’s robustness under more challenging conditions. This configuration serves as an idealized representation of the rank-ordered internal soliton packets widely documented in the South China Sea, the Strait of Gibraltar, and other energetic oceanic environments, where three or more solitary waves of decreasing amplitude propagate as a coherent group [46, 49]. The initial condition generalizes equation (52) to three components:
@@
with amplitudes $A_{1}=7.0$ m, $A_{2}=4.0$ m, $A_{3}=2.5$ m, widths $w_{1}=1.2$ m, $w_{2}=1.8$ m, $w_{3}=2.2$ m, and initial positions $x_{1}=-25.0$ m, $x_{2}=-10.0$ m, $x_{3}=5.0$ m, respectively. The corresponding velocities are $v_{1}\approx 0.467$ m/s, $v_{2}\approx 0.267$ m/s, and $v_{3}\approx 0.167$ m/s, ordered such that each faster soliton will sequentially overtake the slower ones ahead. To accommodate the increased spatial extent of the three-wave system and provide adequate resolution for the narrowest soliton, the domain is enlarged to $x\in[-50,50]$ m with enhanced resolution $N=1024$ ($\Delta x\approx 0.098$ m). The simulation extends to $T_{\text{final}}=80.0$ s with tightened tolerances $\text{rtol}=10^{-11}$ and $\text{atol}=10^{-13}$ to maintain accuracy through multiple collision events. This test case may exhibit complex interaction patterns including transient amplitude amplifications during collisions, the generation of small-amplitude dispersive radiation, and intricate phase shift accumulations [6, 1]. The successful navigation of these multiple interactions while preserving conservation laws and soliton identities demonstrates the sangkuriang solver’s capability for research-grade simulations of multi-soliton dynamics relevant to oceanic internal wave studies.
@@
The physical parameters $\mu=0.1$ m3/s and $\varepsilon=0.2$ m-1 are held constant across all test cases to facilitate direct comparison of dynamical behaviors. These values are representative of the weakly nonlinear, weakly dispersive regime characteristic of long gravity waves in shallow coastal waters [26], and ensure that the balance between nonlinearity and dispersion falls within the regime where soliton solutions are well-defined. The specific choices of soliton amplitudes and positions in each test case are informed by established benchmarks in the soliton literature [3, 20] and are designed to produce qualitatively distinct interaction scenarios within computationally tractable simulation times. While these four cases provide comprehensive validation of the sangkuriang solver’s core functionality, the implementation supports arbitrary initial conditions through the SechProfile, TanhProfile, GaussianProfile, and MultiSoliton classes, enabling users to investigate alternative scenarios such as soliton generation from non-sech2 initial data, wave packet decomposition, or soliton interactions with continuous wave backgrounds—all of which are relevant to understanding observed oceanic internal wave phenomena [50, 51].
@@
The output from each sangkuriang simulation includes the complete spatiotemporal field $u(x,t)$, conservation law time series $M(t)$, $P(t)$, $E(t)$ as defined in equations (45)–(47), relative conservation errors (48)–(50), and diagnostic information regarding computational cost (number of function evaluations, wall clock time). Data are stored in netCDF4 format [23] following CF-1.8 metadata conventions, facilitating interoperability with standard oceanographic data analysis tools and ensuring compatibility with widely used geoscience software packages. Animated visualizations are generated showing the evolving wave profile in three-dimensional space-time representations, with conservation law diagnostics overlaid to provide immediate visual feedback on solution quality. These outputs enable both qualitative assessment of soliton dynamics and quantitative verification of the numerical implementation’s fidelity to the underlying physics encoded in the KdV equation (29).
@@ 2.4 Data Analysis
The diagnostic analysis of simulation outputs is performed independently of the sangkuriang solver library through a dedicated suite of post-processing routines implemented in Python. These routines employ NumPy [11] for array operations and FFT, SciPy [12] for signal processing and distance computations, and Matplotlib [35] for visualization. Simulation data archived in netCDF4 format [23] following CF-1.8 conventions are subjected to four complementary analyses: verification of conservation law preservation, soliton trajectory tracking with velocity validation, spectral information-theoretic characterization, and phase space recurrence quantification. Each analysis extracts quantitative metrics that collectively assess both numerical accuracy and physical fidelity of the computed solutions. Together, these diagnostics provide a comprehensive assessment framework applicable not only to the idealized KdV system studied here but also to more complex ocean wave models where analytical benchmarks may be unavailable [50, 51].
@@
provides a time-averaged measure of conservation fidelity that is less sensitive to isolated outliers. To distinguish systematic drift from bounded oscillations—a distinction important for assessing whether numerical errors would accumulate to unphysical levels over the long propagation distances typical of oceanic solitary waves—the linear drift rate $\gamma_{Q}$ is extracted via ordinary least-squares regression. Defining the design matrix $\mathbf{A}\in\mathbb{R}^{N_{t}\times 2}$ with elements $A_{n1}=t_{n}$ and $A_{n2}=1$, and the observation vector $\mathbf{b}=[\delta_{Q}(t_{0}),\delta_{Q}(t_{1}),\ldots,\delta_{Q}(t_{N_{t}-1})]^{T}$, the drift rate is obtained as the first component of the least-squares solution
@@
Soliton trajectory extraction proceeds through systematic identification and tracking of local amplitude maxima in the spatiotemporal field $u(x_{j},t_{n})$. This procedure is analogous to the peak-tracking methods employed in observational oceanography to identify and follow individual internal solitary waves in mooring time series or satellite image sequences [47, 49]. At each output time $t_{n}$, the discrete spatial profile $\{u_{j}^{(n)}\}_{j=0}^{N-1}\equiv\{u(x_{j},t_{n})\}_{j=0}^{N-1}$ is analyzed using the scipy.signal.find_peaks routine [12], which identifies indices $j^{*}$ satisfying $u_{j^{*}}^{(n)}>u_{j^{*}-1}^{(n)}$, $u_{j^{*}}^{(n)}>u_{j^{*}+1}^{(n)}$, subject to additional constraints. Two filtering parameters govern peak selection: a height threshold $u_{\text{th}}$ requiring $u_{j^{*}}^{(n)}>u_{\text{th}}$, and a minimum separation distance $d_{\text{min}}$ (in grid points) enforcing $|j_{1}^{*}-j_{2}^{*}|>d_{\text{min}}$ between distinct peaks. For the test cases examined, $u_{\text{th}}=0.5$ m for multi-soliton configurations and $u_{\text{th}}=1.0$ m for single-soliton runs, with $d_{\text{min}}=10$ grid points throughout. Each detected peak yields a triplet $(t_{n},x_{j^{*}},u_{j^{*}}^{(n)})$ recording the time, spatial position, and amplitude.
For configurations containing $K$ solitons, individual trajectories must be disambiguated from the aggregate peak data. This is accomplished through amplitude-based sorting at each time instant: letting $\{(x_{k}^{(n)},A_{k}^{(n)})\}_{k=1}^{K_{n}}$ denote the $K_{n}$ peaks detected at time $t_{n}$ with positions $x_{k}^{(n)}$ and amplitudes $A_{k}^{(n)}$, these are reordered such that $A_{1}^{(n)}\geq A_{2}^{(n)}\geq\cdots\geq A_{K_{n}}^{(n)}$. The $k$-th soliton track is then defined as the sequence $\{(t_{n},x_{k}^{(n)})\}$ across all times where at least $k$ peaks were detected. This ranking scheme exploits the amplitude-velocity ordering inherent to KdV solitons: since taller solitons propagate faster according to equation (34), amplitude rank provides a consistent identifier even through collision events where spatial ordering temporarily inverts. The same amplitude-rank ordering is a well-known feature of oceanic internal solitary wave packets, where the leading wave is typically the largest [46, 48].
@@
where terms with $p_{m}^{(n)}<10^{-15}$ are excluded from the summation to avoid numerical singularities. The normalization by $\ln N$ ensures $S_{k}^{(n)}\in[0,1]$, with $S_{k}^{(n)}=0$ corresponding to a monochromatic signal (all energy concentrated at a single wavenumber) and $S_{k}^{(n)}=1$ to uniform spectral distribution (white noise). Soliton profiles, characterized by smooth $\text{sech}^{2}$ envelopes with exponentially decaying Fourier spectra, exhibit intermediate entropy values that remain approximately constant during propagation. Changes in spectral entropy during soliton interactions provide a measure of how the wave field’s spectral complexity evolves—information that is relevant to understanding energy cascades in oceanic internal wave fields [51].
@@
measures the squared Euclidean distance between the observed distribution $\{p_{m}^{(n)}\}$ and the uniform distribution $\{1/N\}$. This construction ensures $C_{\text{LMC}}^{(n)}=0$ for both perfectly ordered states ($H^{(n)}=0$, all energy at one mode) and maximally disordered states ($D^{(n)}=0$, uniform distribution), attaining nonzero values only for intermediate configurations exhibiting nontrivial structure [37]. The LMC complexity thus distinguishes structured nonequilibrium states from both trivial order and featureless randomness, offering a single scalar diagnostic for the degree of organized wave structure in the computed solutions.
@@
with one-sided differences at domain boundaries. Higher Fisher information indicates sharper spatial gradients and hence more localized wave structures; for soliton solutions, $F^{(n)}$ scales inversely with the characteristic width parameter $w$ appearing in equation (33). During soliton collisions, transient increases in $F^{(n)}$ reflect the temporary steepening of gradients as wave profiles overlap, providing a quantitative signature of the interaction events that parallels the steepening observed in field measurements of colliding internal solitary waves [47].
@@
measures the fraction of recurrence points forming diagonal line structures of length at least $\ell_{\min}$ (typically $\ell_{\min}=2$). Here $P(\ell)$ denotes the number of diagonal lines of exactly length $\ell$ in the recurrence matrix, where a diagonal line at offset $k>0$ consists of consecutive entries $R_{i,i+k}=R_{i+1,i+1+k}=\cdots=1$. Diagonal structures arise when segments of the trajectory evolve in parallel, a hallmark of deterministic dynamics on regular (non-chaotic) attractors [42]. For integrable systems such as the KdV equation, DET $\to 1$ reflects confinement to invariant tori, whereas chaotic systems exhibit DET $\ll 1$ due to sensitive dependence on initial conditions disrupting parallel trajectory segments. In the context of ocean wave dynamics, high determinism values provide confidence that the numerical solver faithfully preserves the regular dynamical structure expected of the idealized KdV system, which is a prerequisite for reliable long-time simulations of solitary wave propagation and interaction [6].
@@
3 Results
The sangkuriang solver was validated through four test cases of increasing complexity, executed on a laptop-class workstation (Lenovo ThinkPad P52s with Intel Core i7-8550U processor at 4.0 GHz, 8 logical cores) running Fedora Linux 39 (kernel 6.11.9-100.fc39.x86_64). This hardware configuration represents a modest computational platform typical of those available to individual researchers in oceanography and geoscience departments, permitting assessment of the solver’s practical utility for idealized ocean wave studies without access to high-performance computing resources.
Table 1 summarizes the computational cost for each test case. The single-soliton simulation (Case 1) required 208,202 adaptive time steps to integrate over the 50 s interval, completing in approximately 152 s of wall-clock time. The adaptive time stepper selected finer temporal resolution for this case due to the smaller spatial domain ($L=60$ m) and correspondingly higher wavenumber content relative to Cases 2 and 3. The two equal-amplitude soliton simulation (Case 2) completed in 75.7 s using 98,612 steps, while the overtaking collision (Case 3) required 115,658 steps and 87.6 s. The three-soliton configuration (Case 4), employing doubled spatial resolution ($N=1024$) and tightened tolerances ($\text{rtol}=10^{-11}$, $\text{atol}=10^{-13}$), demanded 551,858 steps and 534 s of computation time. These timings suggest that the Numba-accelerated implementation achieves reasonable throughput for exploratory research into idealized ocean wave dynamics, though the scaling with grid resolution and tolerance warrants consideration for production simulations involving extended ocean domains.